OpenStudy (anonymous):

How to check the converges of this .. (n^3-2n+1/n) / (7n^5 +2sin(n)-1 ) Please help me.

OpenStudy (anonymous):

convergence as \[n \rightarrow \infty\]?

OpenStudy (anonymous):

Hey , Rlsadiz, I need a method to do that, Help me

OpenStudy (anonymous):

use limit evaluation as it approaches infinity\[\lim_{n \rightarrow \infty} = \frac{\frac{n^{3} - 2n + 1}{n}}{7n^{5} + 2\sin(n) - 1}\] is this the correct question?

OpenStudy (anonymous):

rlsadiz, no the question u typed is wrong.the upper part is n^3-2n+(1/n)

OpenStudy (anonymous):

Take out the highest power of n from the numerator and denominator to get (n^3(1-2/n^2+1/n^4))/(n^5(7+2sin(n)/n^5-1/n^5)) You can then ignore anything with an n on the denominator as it approaches zero as n approaches infinity. This leaves you with n^3/7n^5 which equals 1/7n^2, which clearly converges. Hope this is right.

OpenStudy (anonymous):

Did you use any tests, eg, Ratio Test or Limit comparison test ..?

OpenStudy (anonymous):

That test is for series convergence.

OpenStudy (anonymous):

rlsadiz, how do you prove this series is converges then?

OpenStudy (anonymous):

What dr25161 did is right.

OpenStudy (anonymous):

wait are you asking for \[\sum_{x = 1}^{\infty} the function\]?

OpenStudy (anonymous):

rlsadiz, you are right..

OpenStudy (anonymous):

I dont know what that test is called ( i learnt calculus in a foreign language) but its the one where you do A(n+1)/A(n) and if its less than 1, the series converges.

OpenStudy (anonymous):

dr25161, it is called Ratio test, But I'm stuck in the middle way, Please can you solve it step by step.

OpenStudy (anonymous):

You could use comparison test. We need to find a function such that it is larger than our function then prove that it will converge. Consider the function: \[1.)\frac{n^{3}}{7n^{5} - 3}\] it is larger than our function because \[-1\le \sin n \le 1\] and \[0\ge-2n+\frac{1}{n}\] We could now use limit comparison test to evaluate no. 1\[\lim_{n \rightarrow \infty} \left|\frac{\frac{n^{3}}{n^{5} - 3}}{\frac{n^{3}}{n^{5}}} \right|\] Evaluating the limit by getting the highest power we get 1. Therefore it converges.

OpenStudy (anonymous):

Thanks a lot man, it was really helpful...